In a walkathon to raise money for charity Alicia walked a certain distance at 5 mi/h and then jogged twice that distance at 8 mi/h. Her total walking and jogging was 2 hours 15 minutes. How many miles was the walkathon?
Answer by CharlesG2(834)

"In a walkathon to raise money for charity Alicia walked a certain distance at 5 mi/h and then jogged twice that distance at 8 mi/h. Her total walking and jogging was 2 hours 15 minutes. How many miles was the walkathon?"
D distance = R rate * T time
let d = distance walked at 5 mph (mph = mi/h)
let 2d = distance jogged at 8 mph
let 3d = total distance in 2.25 hours (15 minutes = 0.25 hours)
let t = time walking
let 2.25 - t = time jogging
d = 5 * t
2d = 8 * (2.25 - t) = 18 - 8t --> d = 9 - 4t (dividing both sides by 2)
5t = 9 - 4t (setting the 2 d's equal and solving for t)
9t = 9
t = 1 hour
if t = 1 hour then d = 5 miles, then 3d = 15 miles
walkathon is 15 miles